find two consecutive whole numbers with a product of 210
Answer by mananth(16949)

leth one number be x
the second number will be x+1 (consecutive number)
the product of the two = 210
x(x+1)=210
X^2 +X-210=0
X^2+15X-14X-210=0
X(X+15)-14(X+15)=0
(X-14)(X+15) =0
X=14 OR -15
THE NUMBERS CAN BE 14 ,15 OR -14 -15